emmathegreat2007:

PLEASE HURRY Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

NottheMonga2087:

so we know pie equals 3.14

Angle:

Interesting question because it wants you to try to "discover" the value of pi. We know that Circumference = pi * (diameter) This means that in (Part A) we can do: Circumference divided by Diameter = pi Circle A: What is 21.98 divided by 7 = ? Circle B: What is 18.84 divided by 6 = ?

Angle:

The two answers to the above comment will be your answers for Part A

emmathegreat2007:

OK

NottheMonga2087:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NottheMonga2087 part a 21.98/7 Pi= 3.14 part b 18.84/6 pi=3.14 remember that \(\color{#0cbb34}{\text{End of Quote}}\)

emmathegreat2007:

ok

NottheMonga2087:

part c pi has a value of 3.14 in all cases

Angle:

We know that Area of a Circle = pi * (radius) * (radius) This means we find the radius using the diameter, then pi = (Area) divided by ( radius x radius ) Circle A: radius = 3.5 What is (38.465) divided by (3.5 x 3.5) = ? be careful with parentheses Circle B: radius = 3 What is (28.26) divided by (3 x 3) = ?

Angle:

The above comment will get you your answers for part B

NottheMonga2087:

part C pi equals 3.14 it will be the same for both circles and is the same for every Circle

NottheMonga2087:

did that answer your question

NottheMonga2087:

i tried i suck at math

emmathegreat2007:

yes it did ThANK YOU both

NottheMonga2087:

your welcome! ;]